# Post-adiabatic waveform-generation framework for asymmetric precessing binaries

### v2

@@ V.2 Gauge invariance
$\mathring{h}^{\prime(2\text{-}\not\chi_{2},\bm{k})}_{\alpha\beta}$ $=\left[\mathring{h}^{(2\text{-}\not\chi_{2})}_{\alpha\beta}+\left\langle\delta\pi_{i}^{(\not\chi_{2})}\right\rangle\frac{\partial}{\partial\mathring{\pi}^{\prime}_{i}}\mathring{h}^{(1)}_{\alpha\beta}\right]e^{-i\langle\Delta\psi_{\bm{k}}\rangle},$ (186)
@@
$h$ $=\sum_{\bm{k}\in\mathbb{Z}^{3}}\Bigl{\{}\varepsilon\mathring{h}^{(1)}_{\bm{k}}(\mathring{\pi}^{\prime}_{i})+\varepsilon^{2}\Bigl{[}\mathring{h}^{(2\text{-}\not\chi_{2})}_{\bm{k}}(\mathring{\varpi}^{\prime}_{i})$
@@ V.3 Relevance of the secondary spin’s precession
[…]
@@
Here $h^{(2)}_{\bm{k},0}=\Bigl{(}h^{(2\text{-}\not\chi_{2})}_{\bm{k}}+\chi_{\parallel}h^{(2\text{-}\chi_{\parallel})}_{\bm{k}}\Bigr{)}e^{-i\bm{\psi_{k}}}$ is the total second-order $\bm{k}$ mode excluding the secondary spin precession.
@@
Now, to compare these terms to the usual $n$PA counting, we note that the form of the evolution equations (95a)–(95c) immediately implies that $\mathring{\psi}_{\bm{k}}$ can be expanded as
@@
in terms of the ‘slow time’ $\varepsilon t$; cf. Eq. (6). Here $\mathring{\psi}^{(0)}_{\bm{k}}/\varepsilon$ is the 0PA term and subsequent orders are $n$PA. Comparing this to Eq. (196), we see that the precession of the secondary spin contributes to the waveform phase at the same order in $\varepsilon$ as a 2PA effect.
@@
[…]
@@ VI Evolution with asymptotic fluxes
$\left\langle\frac{dJ_{\alpha}}{d\tau}\right\rangle_{\!\tau}=-\mathcal{F}\bigl{[}h,\partial_{\vartheta^{\alpha}_{0}}h\bigr{]}+\mathcal{O}(\varepsilon^{2}s^{0},\varepsilon^{3}).$ (201)
Here $(\vartheta^{\alpha},J_{\alpha})$ are action-angle variables for the linearised MPD system in Kerr spacetime, and $\vartheta^{\alpha}_{0}$ is the angle value at the reference time where we wish to compute the average. On the left, $\langle\rangle_{\tau}$ is an average over proper time around the reference time. On the right, the bilinear operator $\mathcal{F}\bigl{[}h,\partial_{\vartheta^{\alpha}_{0}}h\bigr{]}$ is an average flux to future null infinity and down the black hole horizon, constructed from the linear perturbation $h_{\alpha\beta}$ due to a spinning particle. We use $\mathcal{O}(\varepsilon^{2}s^{0})$ to denote spin-independent 1PA terms.
@@
$\left\langle\frac{dJ_{\alpha}}{dt}\right\rangle=-{\cal F}_{\alpha}+\mathcal{O}(\varepsilon^{2}s^{0},\varepsilon^{3}),$ (202)
@@
${\cal F}_{\alpha}\equiv\sum_{\ell\bm{k}}\frac{\varepsilon_{\alpha}}{4\pi\omega_{\bm{k}}^{3}}\left(|\mathcal{Z}^{\text{out}}_{\ell\bm{k}}|^{2}+\frac{\omega_{\bm{k}}}{\mathcal{P}_{\bm{k}}}|\mathcal{Z}^{\text{down}}_{\ell\bm{k}}|^{2}\right).$ (203)
Here
$\omega_{\bm{k}}\equiv k_{i}\Omega^{i}=k_{i}\Bigl{(}\Omega^{i}_{(0)}+\varepsilon\Omega^{i}_{(1\text{-}\chi_{2})}\Bigr{)},$ (204)
@@
$\left\langle\frac{dP_{i}}{dt}\right\rangle=-\frac{\partial P_{i}}{\partial J_{\alpha}}\mathcal{F}_{\alpha}+\mathcal{O}(\varepsilon^{2}s^{0},\varepsilon^{3}),$ (206)
where it is understood that the Jacobian $\partial P_{i}/\partial J_{\alpha}$ is calculated for a test particle to linear order in spin. The form (206) requires the relationships $P_{i}(J_{\alpha})$, which were recently calculated in closed form by Witzany and collaborators [72]. In particular, the linear spin corrections to the Jacobian $\partial P_{i}/\partial J_{\alpha}$ can be obtained analytically from Eqs. (B.9) and (B.15) of Ref. [72].
@@ VI.1.1 Actions
recall that $\chi_{\perp}$ cannot contribute because the frequency involves an average over the precession phase. Both this and Eq. (210) are expansions in $\varepsilon$ at fixed $P$. Since $P_{i}=\langle P_{i}\rangle$ for a spinning test particle, we can replace $P_{i}$ with $\langle P_{i}\rangle$ in the above expressions. The expressions then also hold true in the presence of self-force in the fixed-constants-of-motion gauge discussed in Sec. IV.4; otherwise, they omit $\mathcal{O}(\varepsilon s^{0})$ self-force terms.
@@ VI.1.2 Averages
Before mapping quantities in the regular expansion onto those in the multiscale expansion, we first convert Eq. (201) to an average over $t$. At the same time, we effectively convert to the formalism in the body of this paper: a 3+1 split in which $t$ is a parameter along trajectories rather than a coordinate. This is achieved by observing that if $J_{\alpha}$ is conjugate to $\vartheta^{\alpha}$ for a Hamiltonian $H$ that generates proper-time evolution [i.e., $d\vartheta^{\alpha}/d\tau=\partial H/\partial J_{\alpha}=\nu^{\alpha}(J_{\beta})$], then $J_{\alpha}$ is also conjugate to $\mathring{\psi}^{\alpha}$ for a Hamiltonian ${\cal H}$ that generates coordinate-time evolution [i.e., $d\mathring{\psi}^{i}/dt=\partial{\cal H}/\partial J_{i}=\Omega^{i}(J_{k})$ and $\mathring{\psi}^{t}=t$[132, 133, 134]. Using ${\cal H}$ in place of $H$ in Grant’s derivation yields
$\left\langle\frac{dJ_{\alpha}}{dt}\right\rangle_{\!t}=-\mathcal{F}\Bigl{[}h,\partial_{\mathring{\psi}^{\alpha}_{0}}h\Bigr{]}\equiv-\mathcal{F}_{\alpha},$ (213)
where $\mathring{\psi}^{t}_{0}\equiv t_{0}$ and $\mathring{\psi}^{i}_{0}\equiv\mathring{\psi}^{i}(t_{0})$. [^13] When using $t$ as a time parameter, Hamilton’s equations only apply for $\alpha=i$. As a consequence, Grant’s derivation only yields Eq. (213) for $\alpha=i$. However, the equation for $\alpha=t$ follows from the others as an equation for the Hamiltonian itself: $\left\langle\frac{d{\cal H}}{dt}\right\rangle=\left\langle\frac{\partial{\cal H}}{\partial J_{i}}\frac{dJ_{i}}{dt}\right\rangle=\left\langle\Omega^{i}\frac{dJ_{i}}{dt}\right\rangle=-{\cal F}\Bigl{[}h,\Omega^{i}\partial_{\mathring{\psi}^{i}_{0}}h\Bigr{]}$. This agrees with the $\alpha=t$ component of Eq. (213) because $J_{t}=-{\cal H}$ and, by virtue of Eq. (215), $h_{\alpha\beta}$ depends on $\mathring{\psi}^{\alpha}_{0}$ in the combination $(\mathring{\psi}^{i}_{0}-\Omega^{i}t_{0})$. Alternatively, we can obtain this result directly from Eq. (201) using the general relation $\langle df/d\tau\rangle_{\tau}=\langle df/dt\rangle_{t}/\langle d\tau/dt\rangle_{t}$ [77, 88] and the Jacobian $\partial\mathring{\psi}^{\beta}/\partial\vartheta^{\alpha}$.
@@
$\mathring{\pi}_{i}(\tilde{t}_{0},\varepsilon)=\mathring{\pi}^{(0)}_{i}(\tilde{t}_{0})+\varepsilon\mathring{\pi}^{(1\text{-}\chi_{2})}_{i}(\tilde{t}_{0})+\mathcal{O}(\varepsilon s^{0});$ (217)
@@
$\Omega^{i}(\tilde{t}_{0},\varepsilon)=\Omega^{i}(\tilde{t}_{0})+\varepsilon\Omega^{i}_{(1\text{-}\chi_{2})}(\tilde{t}_{0})+\mathcal{O}(\varepsilon s^{0}).$ (218)
@@
We similarly consider the flux ${\cal F}_{\alpha}$ on the right-hand side of Eq. (213), which is an integral of products of $h_{\alpha\beta}$ and $\partial_{\mathring{\psi}^{\alpha}_{0}}h_{\alpha\beta}$ over cuts of the horizon and future null infinity, averaged over all time along those surfaces. Here $h_{\alpha\beta}$ is the linear perturbation sourced by a spinning particle on an orbit with constant parameters (217) and corresponding constant frequencies (218), discarding the $\mathcal{O}(\varepsilon s^{0})$ terms in those equations. We can write this metric perturbation as
$h_{\alpha\beta}=\sum_{\bm{k}\in\mathbb{Z}^{3}}h^{(\bm{k})}_{\alpha\beta}\,e^{-i\bigl{(}\mathring{\psi}_{\bm{k}}^{0}+\omega_{\bm{k}}\Delta t\bigr{)}}+\mathcal{O}(\varepsilon^{2}s^{0},\varepsilon^{3},\chi_{\perp})$ (226)
@@
in the notation of previous sections, where $\mathring{\pi}_{i}$ is evaluated at $t_{0}$ and we omit the $\mathcal{O}(\varepsilon s^{0})$ terms in Eqs. (217) and (218). We also omit the term proportional to $\chi_{\perp}$ in the metric perturbation, which can only contribute to the flux at $\mathcal{O}(s^{2})$ due to its oscillatory dependence on the precession phase (just as it could only contribute to the local dynamics at that order). Given Eq. (226), the derivative in $\partial_{\mathring{\psi}^{\alpha}_{0}}h_{\alpha\beta}$ can be replaced by
@@ VI.1.3 Fluxes
The essential point is that $\mathcal{F}_{\alpha}\equiv\mathcal{F}\bigl{[}h,\partial_{\mathring{\psi}^{\alpha}_{0}}h\bigr{]}$ is an identical function of $h_{\mu\nu}$ and $\partial_{\mathring{\psi}^{\alpha}_{0}}h_{\mu\nu}$ regardless of whether $h_{\mu\nu}$ is sourced by a spinning or a nonspinning particle. More concretely, it is an identical function of the mode coefficients $h^{\bm{k}}_{\alpha\beta}$ and frequencies $\omega_{\bm{k}}$ in Eq. (226) regardless of their values. The conversion into Teukolsky amplitudes is likewise independent of the values of $h^{\bm{k}}_{\alpha\beta}$ and frequencies $\omega_{\bm{k}}$. Therefore the expression for $\mathcal{F}_{\alpha}$, and hence for $\langle dJ_{\alpha}/dt\rangle$, in terms of Teukolsky amplitudes and mode frequencies, is functionally the same for a spinning body as for a non-spinning body. It follows that $\mathcal{F}_{\alpha}$ for a spinning particle must be the same function of Teukolsky amplitudes and mode frequencies as Isoyama et al.’s result for $\langle dJ_{\alpha}/dt\rangle$ for a nonspinning particle. This is the result reproduced in Eq. (203).
@@ VI.2 Pragmatic summary
$\frac{d\langle E\rangle}{dt}$ $={\cal F}_{t}+\mathcal{O}(\varepsilon^{2}s^{0}),$ (229a)
$\frac{d\langle L_{z}\rangle}{dt}$ $=-{\cal F}_{\phi}+\mathcal{O}(\varepsilon^{2}s^{0}).$ (229b)
@@ VII.2 Path to a complete 1PA waveform model
At 0PA order, modeling efforts have been aided by a substantial simplification: the forcing functions $F^{(0)}_{i}$ in the evolution equation (3) can be written in terms of asymptotic Teukolsky mode amplitudes at infinity and the primary black hole’s horizon [95]. This means all necessary inputs for a 0PA waveform model ($F^{(0)}_{i}$ and the waveform amplitudes $h^{(1)}_{\bm{k}}$) can be determined directly from the solution to the Teukolsky equation with a point-mass source, avoiding the need to reconstruct the complete first-order metric perturbation or to calculate the complete self-force it exerts [2]. As we have highlighted, the same shortcut is now possible for the secondary spin’s contribution to 1PA waveform models. We expect the same to also hold for the $\chi_{2}$-independent sector of the 1PA dynamics, but such a result has not yet been established. As 1PA waveform models are extended to include spin precession and eccentricity, transient orbital resonances between $\Omega_{r}$ and $\Omega_{\theta}$ will also require careful treatment [149, 76]. There has been considerable progress to that end [4, 150, 151, 152].
@@
Finally we remark that, even before accurate 1PA SF models are extended across the entire parameter space (and into the merger-ringdown regime), their intermediate results may be used in calibrating effective models with more extensive coverage. Such calibration has a long history [2], with Refs. [24, 25, 159] standing as recent examples. The results presented in Refs. [90, 91, 92] could be used to further calibrate these models.

### v3

@@ V.2 Gauge invariance
$\mathring{h}^{\prime(2\text{-}\bcancel{\chi_{2}},\bm{k})}_{\alpha\beta}$ $=\left[\mathring{h}^{(2\text{-}\bcancel{\chi_{2}})}_{\alpha\beta}+\left\langle\delta\pi_{i}^{(\bcancel{\chi_{2}})}\right\rangle\frac{\partial}{\partial\mathring{\pi}^{\prime}_{i}}\mathring{h}^{(1)}_{\alpha\beta}\right]e^{-i\langle\Delta\psi_{\bm{k}}\rangle},$ (186)
@@
$h$ $=\sum_{\bm{k}\in\mathbb{Z}^{3}}\Bigl\{\varepsilon\mathring{h}^{(1)}_{\bm{k}}(\mathring{\pi}^{\prime}_{i})+\varepsilon^{2}\Bigl[\mathring{h}^{(2\text{-}\bcancel{\chi_{2}})}_{\bm{k}}(\mathring{\varpi}^{\prime}_{i})$
@@ V.3 Relevance of the secondary spin’s precession
$\Phi_{\bm{k}}=\mathring{\psi}_{\bm{k}}-\varepsilon^{0}\arg\bigl(h^{(1)}_{\bm{k}}\bigr)+\frac{\varepsilon}{\bigl|h^{(1)}_{\bm{k}}\bigr|^{2}}\Bigl\{d^{\,0}_{\bm{k}}+\chi_{\perp}d^{\,\rm s}_{\bm{k}}\sin\bigl(\mathring{\psi}_{s}\bigr)\\ +\chi_{\perp}d^{\,\rm c}_{\bm{k}}\cos\bigl(\mathring{\psi}_{s}\bigr)\Bigr\}+{\cal O}(\varepsilon^{2}),$ (196)
@@
Here $h^{(2)}_{\bm{k},0}=\Bigl(h^{(2\text{-}\bcancel{\chi_{2}})}_{\bm{k}}+\chi_{\parallel}h^{(2\text{-}\chi_{\parallel})}_{\bm{k}}\Bigr)e^{-i\bm{\psi_{k}}}$ is the total second-order $\bm{k}$ mode excluding the secondary spin precession.
@@
Now, to compare these terms to the usual nnPA counting, we note that the form of the evolution equations (95a)–(95c) immediately implies that $\mathring{\psi}_{\bm{k}}$ can be expanded as
@@
in terms of the ‘slow time’ $\varepsilon t$; cf. Eq. (6). Here $\mathring{\psi}^{(0)}_{\bm{k}}/\varepsilon$ is the 0PA term and subsequent orders are nnPA. Comparing this to Eq. (196), we see that the precession of the secondary spin contributes to the waveform phase at the same order in $\varepsilon$ as a 2PA effect.
@@
However, both of these conclusions might be too hasty. About the first, we note that the effect of the spin precession is qualitatively different than the effect of a 2PA forcing function: unlike a 2PA term, the precession terms (in both the phase and the amplitude) oscillate on a fast time scale rather than only varying on the slow, radiation-reaction time scale of the inspiral. About the second conclusion, we note that first-order conservative and second-order dissipative effects have the same qualitative effect on the waveform, and the division between them is gauge dependent. There is hence no reason to expect that including just one of them will improve the fidelity of a waveform model. This contrasts with the secondary spin precession, which has a qualitatively different signature than other effects that contribute to the waveform phase at the same order in $\varepsilon$. Ultimately, data analysis studies should assess the relevance of the secondary spin precession for gravitational-wave science. In Ref. [92], we highlight that $a^{\,\rm c}_{\bm{k}}$, $d^{\,\rm c}_{\bm{k}}$, $a^{\,\rm s}_{\bm{k}}$ and $d^{\,\rm s}_{\bm{k}}$ are numerically small compared to $a^{\,0}_{\bm{k}}$ and $d^{\,0}_{\bm{k}}$ for quasicircular, approximately equatorial inspirals into slowly spinning primaries. Their relative contribution to the waveform for generic configurations warrants further investigation.
@@ VI Evolution with asymptotic fluxes
$\left\langle\frac{dJ_{\alpha}}{d\tau}\right\rangle_{\!\tau}=-\left\langle\mathcal{F}\bigl[h,\partial_{\vartheta^{\alpha}_{0}}h\bigr]\right\rangle_{\tau}+\mathcal{O}(\varepsilon^{2}\bcancel{s},\varepsilon^{3}).$ (201)
Here we use $\mathcal{O}(\varepsilon^{2}\bcancel{s})$ to denote spin-independent 1PA terms. The variables $(\vartheta^{\alpha},J_{\alpha})$ are action-angle variables for the linearised MPD system in Kerr spacetime, and $\vartheta^{\alpha}_{0}$ is the angle value at the reference time where we wish to compute the average. On the left, $\langle\rangle_{\tau}$ is an average over proper time around the reference time. On the right, the bilinear operator $\left\langle\mathcal{F}\bigl[h,\partial_{\vartheta^{\alpha}_{0}}h\bigr]\right\rangle_{\tau}$ is a certain $\tau$-averaged flux to future null infinity and down the black hole horizon, constructed from the linear perturbation $h_{\alpha\beta}$ due to a spinning particle.
@@
$\left\langle\frac{dJ_{\alpha}}{dt}\right\rangle=-{\cal F}_{\alpha}+\mathcal{O}(\varepsilon^{2}\bcancel{s},\varepsilon^{3}),$ (202)
where $\langle\cdot\rangle$ is our angle average, and (adapting the notation of Ref. [90]) the flux is [^11] For comparison with Ref. [90], we note that the mode numbers there are written as $k_{i}=(n,k,m)$.
${\cal F}_{\alpha}=\sum_{\ell\bm{k}}\frac{\varepsilon_{\alpha}}{4\pi\omega_{\bm{k}}^{3}}\left(|\mathcal{Z}^{\text{out}}_{\ell\bm{k}}|^{2}+\frac{\omega_{\bm{k}}}{\mathcal{P}_{\bm{k}}}|\mathcal{Z}^{\text{down}}_{\ell\bm{k}}|^{2}\right).$ (203)
Here
@@
$\left\langle\frac{dP_{i}}{dt}\right\rangle=-\frac{\partial P_{i}}{\partial J_{\alpha}}\mathcal{F}_{\alpha}+\mathcal{O}(\varepsilon^{2}\bcancel{s},\varepsilon^{3}),$ (206)
where it is understood that the Jacobian $\partial P_{i}/\partial J_{\alpha}$ is calculated for a test particle to linear order in spin. The form (206) requires the relationships $P_{i}(J_{\alpha})$, whose inverse relations $J_{\alpha}(P_{i})$ were recently calculated in closed form by Witzany and collaborators [73]. In particular, the linear spin corrections to the Jacobian $\partial J_{\alpha}/\partial P_{i}$ can be obtained analytically from Eqs. (B.9) and (B.15) of Ref. [73] (and $\partial P_{i}/\partial J_{\alpha}$ from its matrix inverse).
@@ VI.1.1 Actions
recall that $\chi_{\perp}$ cannot contribute because the frequency involves an average over the precession phase. Both this and Eq. (210) are expansions in $\varepsilon$ at fixed $P$. Since $P_{i}=\langle P_{i}\rangle$ for a spinning test particle, we can replace $P_{i}$ with $\langle P_{i}\rangle$ in the above expressions. The expressions then also hold true in the presence of self-force in the fixed-constants-of-motion gauge discussed in Sec. IV.4; otherwise, they omit $\mathcal{O}(\varepsilon\bcancel{s})$ self-force terms.
@@ VI.1.2 Averages
[…]
$\left\langle\frac{dJ_{\alpha}}{dt}\right\rangle_{\!t}=-\left\langle\mathcal{F}\Bigl[h,\partial_{\mathring{\psi}^{\alpha}_{0}}h\Bigr]\right\rangle_{t}\equiv-\mathcal{F}_{\alpha},$ (213)
[…]
@@
$\mathring{\pi}_{i}(\tilde{t}_{0},\varepsilon)=\mathring{\pi}^{(0)}_{i}(\tilde{t}_{0})+\varepsilon\mathring{\pi}^{(1\text{-}\chi_{2})}_{i}(\tilde{t}_{0})+\mathcal{O}(\varepsilon\bcancel{s});$ (217)
@@
$\Omega^{i}(\tilde{t}_{0},\varepsilon)=\Omega^{i}(\tilde{t}_{0})+\varepsilon\Omega^{i}_{(1\text{-}\chi_{2})}(\tilde{t}_{0})+\mathcal{O}(\varepsilon\bcancel{s}).$ (218)
@@
We similarly consider the flux ${\cal F}_{\alpha}$ on the right-hand side of Eq. (213), which is an integral of products of $h_{\alpha\beta}$ and $\partial_{\mathring{\psi}^{\alpha}_{0}}h_{\alpha\beta}$ over cuts of the horizon and future null infinity, averaged over all time along those surfaces. Here $h_{\alpha\beta}$ is the linear perturbation sourced by a spinning particle on an orbit with constant parameters (217) and corresponding constant frequencies (218), discarding the $\mathcal{O}(\varepsilon\bcancel{s})$ terms in those equations. We can write this metric perturbation as
$h_{\alpha\beta}=\sum_{\bm{k}\in\mathbb{Z}^{3}}h^{(\bm{k})}_{\alpha\beta}\,e^{-i\bigl(\mathring{\psi}_{\bm{k}}^{0}+\omega_{\bm{k}}\Delta t\bigr)}+\mathcal{O}(\varepsilon^{2}\bcancel{s},\varepsilon^{3},\chi_{\perp})$ (226)
@@
in the notation of previous sections, where $\mathring{\pi}_{i}$ is evaluated at $t_{0}$ and we omit the $\mathcal{O}(\varepsilon\bcancel{s})$ terms in Eqs. (217) and (218). We also omit the term proportional to $\chi_{\perp}$ in the metric perturbation, which can only contribute to the flux at $\mathcal{O}(s^{2})$ due to its oscillatory dependence on the precession phase (just as it could only contribute to the local dynamics at that order). Given Eq. (226), the derivative in $\partial_{\mathring{\psi}^{\alpha}_{0}}h_{\alpha\beta}$ can be replaced by
@@ VI.1.3 Fluxes
The essential point is that $\mathcal{F}_{\alpha}\equiv\mathcal{F}\bigl[h,\partial_{\mathring{\psi}^{\alpha}_{0}}h\bigr]$ is an identical function of $h_{\mu\nu}$ and $\partial_{\mathring{\psi}^{\alpha}_{0}}h_{\mu\nu}$ regardless of whether $h_{\mu\nu}$ is sourced by a spinning or a nonspinning particle. More concretely, it is an identical function of the mode coefficients $h^{\bm{k}}_{\alpha\beta}$ and frequencies $\omega_{\bm{k}}$ in Eq. (226) regardless of their values. The conversion into Teukolsky amplitudes is likewise independent of the values of $h^{\bm{k}}_{\alpha\beta}$ and frequencies $\omega_{\bm{k}}$. Therefore the expression for $\mathcal{F}_{\alpha}$, and hence for $\langle dJ_{\alpha}/dt\rangle$, in terms of Teukolsky amplitudes and mode frequencies, is functionally the same for a spinning body as for a non-spinning body. It follows that $\mathcal{F}_{\alpha}$ for a spinning particle must be the same function of Teukolsky amplitudes and mode frequencies as Isoyama et al.’s result for $\langle dJ_{\alpha}/dt\rangle$ for a nonspinning particle. This is the formula reproduced in Eq. (203).
@@ VI.2 Pragmatic summary
$\frac{d\langle E\rangle}{dt}$ $={\cal F}_{t}+\mathcal{O}(\varepsilon^{2}\bcancel{s}),$ (229a)
$\frac{d\langle L_{z}\rangle}{dt}$ $=-{\cal F}_{\phi}+\mathcal{O}(\varepsilon^{2}\bcancel{s}).$ (229b)
@@ VII.2 Path to a complete 1PA waveform model
[…]
@@
Finally we remark that, even before accurate 1PA SF models are extended across the entire parameter space (and into the merger-ringdown regime), their intermediate results may be used in calibrating effective models with more extensive coverage. Such calibration has a long history [2], with Refs. [25, 26, 161, 162] standing as recent examples. The results presented in Refs. [92, 93, 94] could be used to further calibrate these models.